Circular No. 8321

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 DZ_{61} (CATALINA-LINEAR) An apparently asteroidal object reported independently by the Cata-

lina and LINEAR surveys (discovery observations below from MPEC 2004-E02 and MPS 100749) has been found to have a 4′′-diameter centrally condensed coma and a faint tail in p.a. 190^{o}-220^{o} on R-band CCD images taken on Mar. 16 and 17 UT by S. S. Sheppard, Y. R. Fernandez, and D. Jewitt with the University of Hawaii 2.2-m reflector at Mauna Kea.

2004 UT α_{2000} δ_{2000} Mag. Observer
Feb. 18.45275 12^{h}37^{m}20.^{s}26 −1^{o}52′38′′.3 19.3 Catalina
19.37483 12 37 19.80 −1 19 43.6 19.3 LINEAR

Additional astrometry, the following elliptical orbital elements (P = 308 yr), and an ephemeris appear on MPEC 2004-G34.

Epoch = 2004 June 4.0 TT

T = 2004 May 26.8576 TT  ω = 44.^{o}4612
e = 0.955901  Ω = 172.7922 }2000.0
q = 2.013352 AU  i = 66.8102

SUPERNOVA 2004bi IN UGC 5894 Further to IAUC 8317, K. Shimasaki and W. Li report the LOSS dis-

covery, on KAIT images taken on Apr. 13.23 and 14.20 UT, of an apparent supernova (mag ∼ 18.7) located at α = 10^{h}47^{m}37^{s}.45, δ = +26^{o}18′12′′.0 (equinox 2000.0), which is 25′′.8 west and 32′′.1 north of the nucleus of UGC 5894. A KAIT image taken on Mar. 15.25 showed nothing at this position (limiting mag ∼ 19.8).

SUPERNOVA 2004bg IN UGC 6363 T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2004bg (cf. IAUC 8317), obtained by P. Berlind on Apr. 11.39 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 6306 km/s for the host galaxy, the supernova expansion velocity is 10600 km/s for Si II (rest 635.5 nm). The spectralfeature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 3 ± 2 days before maximum light.

2004 April 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green